Determine the boiling points of the following solutions. All are ionic compoun a. 16.0 grams of CaCl2 dissolved in 1000.0 grams of water. CaC2 110.9 g/mol b. 23.0 grams of CusO, dissolved in 1000.0 grams of water. Cuso 159.6 /m c. 30.0 grams of AgNOs dissolved in 1000 grams of water. AgNO 169.9 g/mo
Answer #1

a) DTb = i*kb*m

DTb = Tb-T0

T0 = boiling point of solvent(water) = 100 C

Tb = boiling point of solution = x C

m = molality of solution = (w/M)*(1000/ W in g)

   = (16/110.9)*(1000/1000)

i = vanthaff factor of CaCl2 = 3

kb of water = 0.512 °C/m

(x-100) = 3*0.512*(16/110.9)*(1000/1000)



boiling point of solution = 100.22 c

b) DTb = i*kb*m

DTb = Tb-T0

T0 = boiling point of solvent(water) = 100 C

Tb = boiling point of solution = x C

m = molality of solution = (w/M)*(1000/ W in g)

   = (23/159.6)*(1000/1000)

i = vanthaff factor of CuSO4 = 2

kb of water = 0.512 °C/m

(x-100) = 2*0.512*(23/159.6)*(1000/1000)



boiling point of solution = 100.15 c

C)

DTb = i*kb*m

DTb = Tb-T0

T0 = boiling point of solvent(water) = 100 C

Tb = boiling point of solution = x C

m = molality of solution = (w/M)*(1000/ W in g)

   = (30/169.9)*(1000/1000)

i = vanthaff factor of AgNO3 = 2

kb of water = 0.512 °C/m

(x-100) = 2*0.512*(30/169.9)*(1000/1000)



boiling point of solution = 100.18 c